# Doubly heavy tetraquarks from lattice QCD: incorporating diquark-antidiquark operators and the left-hand cut

### v1

@@ Abstract
Lattice studies of the doubly-charm tetraquark $T_{cc}=cc\bar{u}\bar{d}$ require the determination of the $DD^{*}$ scattering amplitude, which most often incorporate only meson-meson interpolators. We additionally incorporate diquark-antidiquark operators and find that these have some impact on certain eigenenergies, render slightly smaller values of $p\cot\delta_{0}$ for $DD^{*}$ scattering and a $T_{cc}$ pole slightly closer to the threshold. A more significant effect of diquark-antidiquark operators on eigen-energies is found for larger heavy quark masses relevant for $T_{bb}$. The $DD^{*}$ scattering amplitude is extracted from eigenenergies by adopting plane-wave and effective-field-theoretic methods, which also incorporate the left-hand cut and address the partial wave mixing. The $T_{cc}$ is found to be a subthreshold resonance with a pole at $m_{T_{cc}}-m_{D}-m_{D^{*}}=-5.2^{+0.7}_{-0.8}-i\cdot 6.3^{+2.4}_{-4.8}\leavevmode\nobreak$eV. Lattice simulations were performed on CLS ensembles with $m_{\pi}\simeq 280\leavevmode\nobreak$MeV using the distillation method.
@@ II Lattice Setup and heavy-light meson masses
The light and heavy quarks are based on the nonperturbatively $O(a)$ improved Wilson-Clover action. Two values of the heavy quark mass are employed: one quark mass is slightly larger than the physical charm quark mass, and the second one, representing the bottom quark, has a smaller-than-physical $b$-quark mass to avoid too large cutoff effects, which might qualitatively affect the inferences derived here. The relevant heavy-meson masses for both heavy quarks are provided in Table 2.
@@ III Operator basis
Possible effects of local four-quark operators (4) have not been explored extensively in lattice studies at masses closer to that of the charm quark mass. Their effect on eigen-energies in the $T_{cc}$ channel has been found to be insignificant [9, 10] or mild [11, 13, 12]. This is the first work where their effect on the extracted $DD^{*}$ scattering amplitude has been explored. Preliminary results, which showed the local four-quark operators may have a mild effect, were presented in our proceedings [13].
@@ V Finite-volume energies, overlaps, and impact of local four-quark operators
    The observation of this lower level with the enlarged basis indicates that bilocal meson-meson interpolators fail to access the ground state in the $T_{bb}$ sector within moderate physical time separations. With the inclusion of the local diquark-antidiquark operator, we obtain a $T_{bb}$ binding energy of approximately $60\leavevmode\nobreak\ \pm 10$ MeV for the lighter-than-physical "$b"$ quark we investigate (see Table 2). The deep binding of $T_{bb}$ was reported by previous lattice studies [20, 21] that used local diquark-antidiquark operators alongside local and bilocal meson-meson interpolators in their analysis. Several other studies also have reported similar deep binding using only local $BB^{*}=(b\bar{d})(b\bar{u})$ operators together with local diquark-antidiquark interpolators [19, 10, 40, 41, 22] [^3] Lattice investigations following static potential-based and HALQCD-potential based approaches also predict similar binding energies in $T_{bb}$ system [36, 37, 42]. The only previous lattice simulation of $T_{bb}$ that evaluates all correlators between bi-local $B(0)B^{*}(0)$ and local $[bb][\bar{u}\bar{d}]$ did not present the result based on bilocal $B(0)B^{*}(0)$ operators alone [21].
@@ VII.5 Outlook
Various alternative ways of dealing with the left-hand cuts in scattering amplitudes have been proposed. For example, in Ref. [48] the two-particle Lüscher’s formalism is generalized to explicitly account for the left-hand cut by relaxing some requirements of the original quantization condition. Similarly, [49] proposed a modified quantization condition in the presence of long-range forces that usually arise due to exchanges of light particles, such as one-pion exchange explicitly treated in this work. In addition to these, Refs. [50, 51] develop and apply to the $T_{cc}$, respectively, the 3-body quantization condition that effectively extends its range of validity all the way up to the first 4-particle states that can go on shell in a given channel. In this formalism $D^{*}$ features as $D\pi$ bound state when $m_{\pi}>m_{D^{*}}-m_{D}$.
@@ VIII Conclusions
The $DD^{*}$ scattering amplitude was extracted from lattice eigen-energies in a framework combining an effective field theory and plane-wave approach. This framework is applicable also for energies on the left-hand cut, which is present in the lattice kinematics with stable $D^{*}$. Three low-energy constants of the EFT potential were fitted from nine lattice energies leading to a favorable reproduction of the lattice data. The resulting scattering amplitude agrees with the one obtained with the Lüscher’s approach in the energy region above the left-hand cut where the latter is applicable.

### v2

@@ Abstract
Lattice studies of the doubly-charm tetraquark $T_{cc}=cc\bar{u}\bar{d}$ require the determination of the $DD^{*}$ scattering amplitude, which most often incorporate only meson-meson interpolators. We additionally incorporate diquark-antidiquark operators and find that these have some impact on certain eigenenergies. This study presents the first extraction of the $DD^{*}$ scattering amplitude based on the meson-meson as well as diquark antidiquark interpolators. The effect of the additional operators renders slightly smaller values of $p\cot\delta_{0}$ and a $T_{cc}$ pole slightly closer to the threshold. The scattering amplitude is extracted from eigenenergies by adopting plane-wave and effective-field-theoretic methods, which also incorporate the left-hand cut and address the partial wave mixing. The $T_{cc}$ is found to be a subthreshold resonance with a pole at $m_{T_{cc}}-m_{D}-m_{D^{*}}=-5.2^{+0.7}_{-0.8}-i\cdot 6.3^{+2.4}_{-4.8}\leavevmode\nobreak$eV, employing CLS ensembles with $m_{\pi}\simeq 280\leavevmode\nobreak$MeV and the distillation method. A more significant effect of diquark-antidiquark operators on eigen-energies is found for larger heavy quark masses relevant for $T_{bb}$. We find that deeply bound $T_{bb}$ does not emerge when employing only meson-meson operators, where each meson is separately momentum projected, while the deeply bound state emerges after adding local diquark antidiquark operators.
@@ II Lattice Setup and heavy-light meson masses
The light and heavy quarks are based on the relativistic nonperturbatively $O(a)$ improved Wilson-Clover action with the clover term set equal to $c_{sw}=1.986246$ for all quarks, and different quark masses realized by different hopping parameters $\kappa$. Two values of the heavy quark mass are employed: one quark mass is slightly larger than the physical charm quark mass, and the second one, representing the bottom quark, has a smaller-than-physical $b$-quark mass to avoid too large cutoff effects, which might qualitatively affect the inferences derived here. The employed hopping parameters for light, charm and bottom are $\kappa_{u/d}=0.13697$, $\kappa_{c}=0.12315$ and $\kappa_{b}=0.0877$, and the relevant pion and heavy-meson masses are provided in Tables 1 and 2.
@@ III Operator basis
Possible effects of local four-quark operators (4) have not been explored extensively in lattice studies at masses closer to that of the charm quark mass. Their effect on eigen-energies in the $T_{cc}$ channel has been found to be insignificant [9, 10] or mild [11, 13, 12]. The present paper represents the first study where the effect of diquark antidiquark operators on the extracted $DD^{*}$ scattering amplitude is explored. Preliminary results of this study, which showed the local four-quark operators may have a mild effect, were presented in our proceedings [13].
@@ V Finite-volume energies, overlaps, and impact of local four-quark operators
    The observation of this lower level with the enlarged basis indicates that bilocal meson-meson interpolators fail to access the ground state in the $T_{bb}$ sector within moderate physical time separations. With the inclusion of the local diquark-antidiquark operator, we obtain a $T_{bb}$ binding energy of approximately $60\leavevmode\nobreak\ \pm 10$ MeV for the lighter-than-physical "$b"$ quark we investigate (see Table 2). The deep binding of $T_{bb}$ was reported by previous lattice studies [20, 21] that used local diquark-antidiquark operators alongside local and bilocal meson-meson interpolators in their analysis. These two studies employed nonrelativistic QCD for bottom quarks, while our present study confirms this finding with relativistic bottom quarks. Several other studies also have reported similar deep binding using only local $BB^{*}=(b\bar{d})(b\bar{u})$ operators together with local diquark-antidiquark interpolators [19, 10, 40, 41, 22] [^3] Lattice investigations following static potential-based and HALQCD-potential based approaches also predict similar binding energies in $T_{bb}$ system [36, 37, 42]. The only previous lattice simulation of $T_{bb}$ that evaluates all correlators between bi-local $B(0)B^{*}(0)$ and local $[bb][\bar{u}\bar{d}]$ did not present the result based on bilocal $B(0)B^{*}(0)$ operators alone [21]. In addition to the heavy-quark treatment mentioned above, there are further differences between Ref. [21] and our study: the former employed stochastic noise propagators for evaluating Wick contractions and we employ distillation method; the former is based on the staggered action for light quarks and we use Wilson-clover action.
@@ VII.5 Outlook
Various alternative ways of dealing with the left-hand cuts in scattering amplitudes have been proposed. For example, in Ref. [48] the two-particle Lüscher’s formalism is generalized to explicitly account for the left-hand cut by relaxing some requirements of the original quantization condition, while more recently this has been extended to coupled channels and arbitrary spin in [49]. Ref. [50] proposed a modified quantization condition in the presence of long-range forces that usually arise due to exchanges of light particles, such as one-pion exchange explicitly treated in this work, while Ref. [51] provides another strategy to deal with these challenges. In addition to these, Refs. [52, 53] develop and apply to the $T_{cc}$, respectively, the 3-body quantization condition that effectively extends its range of validity all the way up to the first 4-particle states that can go on shell in a given channel. In this formalism $D^{*}$ features as $D\pi$ bound state when $m_{\pi}>m_{D^{*}}-m_{D}$.
@@ VIII Conclusions
Our study presents the first extraction of $DD^{*}$ scattering amplitude based on meson-meson as well as diquark antidiquark operators. The scattering amplitude was extracted from lattice eigen-energies in a framework combining an effective field theory and plane-wave approach. This framework is applicable also for energies on the left-hand cut, which is present in the lattice kinematics with stable $D^{*}$. Three low-energy constants of the EFT potential were fitted from nine lattice energies leading to a favorable reproduction of the lattice data. The resulting scattering amplitude agrees with the one obtained with the Lüscher’s approach in the energy region above the left-hand cut where the latter is applicable.